OpenStudy (anonymous):

"Find the constants m and b in the linear function f(x) = mx + b so that f(2) = 10 and the straight line represented by f has slope -5." For some reason, I can't figure this out. Any help + the process would be greatly appreciated!

OpenStudy (ybarrap):

f(x) = mx + b m = slope = -5 f(2) = 10 = (-5)(2) + b b = 10 + 10 = 20 Ans. m=-5, b = 20